Shift report 20010826 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Aug 26 2001 - 16:07:35 EDT

  • Next message: Flemming Videbaek: "dN/deta plots"

    ------------------------------------------------------------
    Date: Sun Aug 26 16:07:34 2001, Shift: 08:00-16:00
    Supervisor: Andrei Makeev, with: 
    Log book#8 pages 25 - 27
    
    Shift Summary:
    ------------------------------------------------------------
    1.There was a beam for short time, but later QLI have
    occured, so not much useful data were taken.
    2. FV put 2 attenuators & 2 splitters on channels 1 and 2
    on left ZDC (before the run 4438) to investigate noise
    reduction and saturation in ZDCs. Run 4438 was just 8
    minutes long (started 12:08 pm, died 12:16 pm). Since that
    time no beam. Polarity of all the magnets set to "A". No
    data were collected at this setting yet (supposed to take
    2M events with "A" polarity).
    3. No serious detector problems, technician from RHIC group
    replaced the current-monitor for D2 magnet since it was
    faulty. The new one shows the current ~ 5 amps higher then
    in reality.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Aug 26 2001 - 16:08:36 EDT